# AdvReal: Adversarial Patch Generation Framework with Application to Adversarial Safety Evaluation of Object Detection Systems

### v1

@@ 3.2 Adversarial Generation Framework
To address the limitations of existing methods, this paper introduces a novel framework for generating adversarial patches designed to deceive human detectors in real-world scenarios. The proposed framework, named AdvReal, jointly optimizes 2D and 3D adversarial losses, effectively tackling challenges such as high intra-class variability, multi-angle attack scenarios, and complex real-world factors. Unlike traditional adversarial patches applied to traffic signs or rigid materials, our method focuses on patches on non-rigid clothing, ensuring effectiveness from multiple angles in realistic scenarios.
As illustrated in Fig.2, the AdvReal framework consists of three key modules: the Patch Adversarial Module, the Realistic Enhancement Module, and the Joint Optimization Module. These modules work collaboratively to enhance the adversarial performance and realism of the generated patches.
@@ 3.2.1 Patch Adversarial Module
To enhance the generalization performance of intra-class adversarial attacks against human detectors, the Patch Adversarial Module leverages a diverse human dataset for 2D adversarial training. In this module, patches are placed at the center of the human bounding boxes within the dataset, followed by transformation operations such as rotation, resizing, offsetting, and occlusion. These transformations ensure the robustness of the adversarial patches under varying conditions. The synthesized images, comprising patches overlaid on humans and their backgrounds, are then utilized for 2D adversarial training to improve attack efficacy.
3.2.2 Realistic Enhance Module
This module employs 3D meshes to model humans, clothing, and pants, attaching the texture map generated by the adversarial patch to the clothing. With the time-space enhance strategy refer to Sec.3.3, 3D adversarial clothes with reasonable folds are rendered. The realistic matching mechanism refer to Sec.3.4 processes the rendered human images by first aligning them to realistic distance, size, elevation, and azimuth to ensure accurate spatial representation. And then applying nonlinear adjustments to brightness and contrast to enhance the lighting conditions for improved realism, seamlessly integrating both geometric and lighting aspects to produce highly realistic human images. Finally, the resized human images are synthesized onto decomposed background images for 3D adversarial training. Notably, the Patch Adversarial Module and the Realistic Enhance Module operate synchronously during each batch training process.
@@ 3.2.3 Joint Optimization Module
The synthetic images produced by the Patch Adversarial Module and the Realistic Enhance Module are fed into the pre-trained human detector targeted for attack. During this process, the detector randomly freezes certain layers, and the output is used to compute the adversarial loss. The appearance of the adversarial patch is iteratively optimized by minimizing a combined loss function, which includes 2D detection loss, 3D detection loss, and Total Variation (TV) loss. The joint optimization mechanism refer to Sec.3.5 ensures that the disguised patch achieves high effectiveness across both 2D and 3D detection scenarios.
@@ 3.3 Non-rigid surfaces modeling
Methods based on shear transformation or 3D thin plate spline(3D-TPS) disregarding the nonlinear deformation characteristics of real fabrics Wei et al. (2024). Such an simplification gives rise to two crucial shortcomings in adversarial pattern deployment: (1) geometric discontinuities in high-curvature regions of cloth folds; and (2) physically implausible stretching artifacts during dynamic body movements, leading to noticeable texture distortions.
To improve realism in fabric deformation, we integrate a physics-driven approach that analyzes the motion flow of the clothes mesh to estimate load-bearing stress points. These stress points serve as key locations for generating natural, non-rigid folds, ensuring diverse but realistic fabric behavior even under consistent adversarial training motions Gundogdu et al. (2020).
@@ 3.4.1 Time-space Mapping
These derived parameters guide the rendering process, where a differentiable renderer projects the adversarial texture onto a 3D human mesh, adjusting its distance, size, and azimuth. Crucially, the rendered model’s silhouette (Fig. 4 (d)) is constrained to fit within the bounding box $b_{i}$’s coordinates (Fig. 4 (c)), ensuring geometric consistency when composited into the background. Subsequently, the human model $H(k)$ is used for combined rendering with the clothing $C(k)$ after non-rigid surface modeling. And the bounding box $B(k)$ and the background image slice $T(k)$ are used as input in Relighting mapping.
@@ 3.5.1 Detection Loss
The detection loss $L_{det}$ operates by prioritizing bounding box predictions that achieve both high localization accuracy and confident classification, shown in eq. 13.
$\mathcal{L}_{det}(p)=max(s_{i,j}\mid IoU(b_{i,j},b_{gt,j}))\geq\tau,l_{i,j}=1$ (13)
Here, $L_{det}$ evaluates predictions based on spatial alignment and semantic relevance. For each candidate box $b_{i,j}$, the loss ensures geometric consistency by verifying if its IoU with the ground truth $b_{gt,j}$ exceeds a threshold $\tau$. Simultaneously, the binary indicator $l_{i,j}$ confirms its validity as a positive sample. The loss then selects the highest confidence score $s_{i,j}$ from all qualified candidates as the value of detection loss. In this study, both 2D detection loss $L_{det}^{patch}$ and 3D detection loss $L_{det}^{real}$ are calculated using this method.
@@ 3.5.3 Patch Optimize
During adversarial sample training, the detection model is treated solely as a fixed closed-box for loss calculation and gradient propagation during patch optimization. Inspired by stochastic depth Huang et al. (2016, 2023), we employ the shakedrop regularization mechanism that randomly adjusts both the forward-output and backward-gradient flows in the residual block Yamada et al. (2018). Specifically, in the forward pass, we combine the identity mapping with the stacked layers’ output to generate new feature representations. For an input feature $\ x_{\text{in}}$ and the stacked layers’ output $H(x_{\text{in}})$, the fusion rule is formulated as:
@@
[…]
@@
where $f_{h}:[-1,1]^{d}\to\mathbb{R}^{C}$ is a pre-trained human detection model mapping the adversarial image $x^{\text{adv}}$ to class probabilities. The function $g(\cdot)$ renders the adversarial patch onto the 3D human model, ensuring its adaptation to the surface geometry.
@@ 4.1.1 Dataset
In the Patch Adversarial Module (2D adversarial), we employ the INRIA Person dataset Dalal and Triggs (2005) to train and test our adversarial patch, which comprises 462 training images and 100 test images.
In the Realistic Adversarial Module (3D adversarial), we randomly selected 562 traffic scene images from the nuScenes dataset Caesar et al. (2020), captured from six camera perspectives (Front, Front Right, Front Left, Back Right, Back Left, and Back), to serve as backgrounds for 3D rendering in evaluating our adversarial patches. Among these, 450 images were captured under adequate lighting and 112 under poor lighting. The daytime and nighttime images were randomly divided into a training set of 462 images and a test set of 100 images.
@@ 4.1.2 Victim Detector
We perform adversarial training and attacks on current mainstream object detectors, including single-stage, two-stage, and transformer-based architectures (shown in Tab. 1). The models are pre-trained on MS COCO dataset Lin et al. (2014). In particular, we conduct closed-box attack tests on SOTA detectors (YOLO-v8, v11, v12) in the physical world for the first time. Floating Point Operations (FLOPs) quantify the computational workload required by a model during inference and serve as an indicator of the model’s complexity. In general, models with higher complexity tend to exhibit improved robustness and generalization capabilities.
@@
YOLO-v2 Single-stage 5.58 2017 Redmon and Farhadi (2017)
YOLO-v3 193.89 2018 Redmon and Farhadi (2018)
YOLO-v5 7.70 2020 Jocher et al. (2020)
YOLO-v8 Single-stage (SOTA) 8.7 2023 Jocher et al. (2023)
YOLO-v11 6.50 2024 Jocher and Qiu (2024)
YOLO-v12 6.50 2025 Tian et al. (2025)
Faster-RCNN Two-stage 180.00 2016 Ren et al. (2016b)
D-DETR Transformer-based 78.00 2021 Zhu et al. (2020)
@@
Method ASR $\uparrow$ (Glass-box) ASR $\uparrow$ (Closed-box) ASR $\uparrow$ (Closed-box Avg)
@@
4.1.3 Evaluation Metric
In the evaluation framework, we define that if the IoU between the detection box and the true box is less than $0.5$, or the confidence is lower than $0.5$, the attack is considered successful. We use multiple quantitative metrics to comprehensively evaluate the effectiveness of algorithm-generated adversarial patches, as follows:
  • 1.
    Attacking Success Rate (ASR): The percentage of adversarial samples that cause the detector to fail.
  • 2.
    Precision: The ratio of correct detections to total detections. Lower precision indicates that the detector is more effectively misled by the adversarial patch.
  • 3.
    Recall: The ratio of correctly detected ground-truth targets to all ground-truth targets. Lower recall signifies that more targets are missed due to the attack.
  • 4.
    F1-Score: The harmonic mean of precision and recall. A lower F1-score reflects an overall degradation in detection performance.
  • 5.
    Average Confidence (AC): The mean confidence score of the detector’s predictions. A lower AC indicates that the detector’s certainty in its detections is reduced due to the adversarial attack, reflecting weakened reliability in its outputs.
4.1.4 Experimental Details
During the digital train and evaluation, the distance from 3D human model to camera is limited in $1$ to $4$ meters. All rendered human models are at random angles and positions, and follow the principle of time-space mapping in Sec. 4.
4.1.5 Baseline
We compared our method AdvReal with both classic and state-of-the-art (SOTA) adversarial methods. The classic methods include AdvPatchBrown et al. (2017), AdvTshirtXu et al. (2020), and NatPatchHu et al. (2021), while the SOTA methods comprise AdvTextureHu et al. (2022), AdvCaTHu et al. (2023), and T-SEAHuang et al. (2023). In addition, we assess White, Gray, and Random Noise patches as control groups. For fairness, the areas of all adversarial patches on both rendered and real garments are kept consistent across digital and physical experiments. With the exception of AdvCat, which is entirely camouflaged, the remaining patches and their corresponding garments are shown in Fig. 5.
Table 3: ASRs(%) under different IoU thresholds for YOLOv2 and Faster-RCNN detectors
Method YOLOv2 $\uparrow$ Faster RCNN $\uparrow$
@@ 4.1.5 Baseline
4.2 Digital Adversarial Attack Experiment
4.2.1 Comparisons to SOTA Methods
In the glass-box comparison, influenced by realistic digital world samples, the ASR of previous adversarial patches dropped significantly relative to the results reported in their original papers. Moreover, the ASR of the natural-looking methods—NatPatch ($6.06\%$) and AdvCat ($0.00\%$)—is only marginally higher than that of the three non-adversarial textures. We attribute this lower performance to training approaches that do not sufficiently account for realistic human body dimensions and authentic lighting conditions. Under closed-box comparison, AdvReal demonstrates superior performance compared to other adversarial methods, including state-of-the-art approaches such as AdvTexture and T-SEA. Notably, when evaluated against the advanced detector YOLOv12, AdvReal achieves an ASR of $70.13\%$, far exceeding that of T-SEA ($21.65\%$). However, our method’s performance on Faster R-CNN and D-DETR declines, likely due to architectural differences between these detectors and the one used for adversarial training.
To further validate AdvReal’s adversarial performance, we evaluated the ASR of various adversarial patches under different IoU thresholds. As shown in Tab. 3, AdvReal achieves higher ASRs than other attack methods when using both YOLOv2 and Faster R-CNN as detectors. Under the YOLOv2 detector, AdvReal attains an ASR of $45\%$ at an IoU threshold of $0.1$—surpassing T-SEA’s $29.00\%$ ASR at an IoU threshold of $0.5$. With the Faster R-CNN detector, AdvReal’s ASR approaches $100\%$ at an IoU threshold of $0.5$, significantly exceeding T-SEA’s $67.10\%$. These results indicate that our algorithm performs effectively even under low thresholds and can successfully attack nearly all samples when the threshold is high.

### v2

@@ 3.2 Adversarial Generation Framework
To address the limitations of existing methods, this paper introduces a novel framework for generating adversarial patches designed to deceive human detectors in real-world scenarios. The proposed framework, named AdvReal, performs joint optimization over 2D and 3D adversarial losses, effectively tackling challenges such as high intra-class variability, multi-angle attack scenarios, and complex real-world factors. Unlike traditional adversarial patches applied to traffic signs or rigid materials, our method focuses on patches applied to non-rigid clothing surfaces, ensuring effectiveness from multiple angles in realistic scenarios.
As illustrated in Fig. 2, the AdvReal framework consists of three key modules: the Patch Adversarial Module, the Realistic Enhancement Module, and the Joint Optimization Module. These components function in synergy to improve both the adversarial strength and the physical plausibility of the generated patches.
@@ 3.2.1 Patch Adversarial Module
To enhance the generalization performance of intra-class adversarial attacks against human detectors, the Patch Adversarial Module leverages a diverse human dataset for 2D adversarial training. In this module, patches are placed at the center of the human bounding boxes within the dataset, followed by transformation operations such as rotation, resizing, offsetting, and occlusion. These transformations are applied to ensure the robustness of the adversarial patches under varying conditions. The synthesized images, comprising patches overlaid on humans and their backgrounds, are then utilized for 2D adversarial training to improve attack efficacy.
3.2.2 Realistic Enhancement Module
This module employs 3D meshes to model humans, clothing, and pants, attaching the texture map generated by the adversarial patch to the clothing. Using the time-space enhancement strategy (see Sec. 3.3), 3D adversarial garments with realistic folds are rendered. The realistic matching mechanism (see Sec. 3.4) first aligns the rendered human models to realistic spatial configurations (distance, scale, elevation, and azimuth), ensuring accurate spatial representation. Then, nonlinear adjustments to brightness and contrast are applied to simulate diverse lighting conditions, seamlessly integrating geometric and illumination variations to produce highly realistic adversarial images. Finally, the resized human images are synthesized onto decomposed background images for 3D adversarial training. Importantly, the Patch Adversarial Module and the Realistic Enhancement Module operate synchronously within each training batch.
@@ 3.2.3 Joint Optimization Module
The synthetic images produced by the Patch Adversarial Module and the Realistic Enhancement Module are fed into the pre-trained human detector targeted for attack. During this process, certain layers of the detector are randomly frozen, and the output is used to compute the adversarial loss. The appearance of the adversarial patch is iteratively optimized by minimizing a combined loss function, which includes 2D detection loss, 3D detection loss, and Total Variation (TV) loss. The joint optimization mechanism (see Sec. 3.5) ensures that the disguised patch remains effective across both 2D and 3D detection scenarios.
@@ 3.3 Non-rigid surfaces modeling
Methods based on shear transformation or 3D thin plate spline (3D-TPS) often disregard the nonlinear deformation characteristics of real fabrics [39]. The simplification leads to two key limitations in adversarial pattern deployment: (1) geometric discontinuities in high-curvature regions of cloth folds, and (2) physically implausible stretching artifacts during dynamic body movements, resulting in noticeable texture distortions.
To improve realism in fabric deformation, we integrate a physics-based modeling strategy that analyzes the motion flow of the clothing mesh to estimate load-bearing stress points. These stress points are used as anchors for generating natural, non-rigid folds, ensuring diverse yet realistic fabric behavior even under consistent adversarial training motions [41].
@@ 3.4.1 Time-space Mapping
These derived parameters guide the rendering process, where a differentiable renderer projects the adversarial texture onto a 3D human mesh, adjusting its distance, size, and azimuth. Crucially, the rendered model’s silhouette (Fig. 4 (d)) is constrained to fit within the bounding box $b_{i}$’s coordinates (Fig. 4 (c)), ensuring geometric consistency when composited into the background. Subsequently, the human model $H(k)$ is used for combined rendering with the clothing $C(k)$ after non-rigid surfaces modeling. And the bounding box $B(k)$ and the background image slice $T(k)$ are used as input in Relighting mapping.
@@ 3.5.1 Detection Loss
The detection loss $\mathcal{L}_{\text{det}}$ operates by prioritizing bounding box predictions that achieve both high localization accuracy and confident classification, as defined in Eq. 13 and Eq. 14.
$\mathcal{L}_{\text{det}}(p)=\text{Conf}_{i^{*}}(p)$ (13)
where
$i^{*}=\text{arg}\max(s_{i,j}\mid\text{IoU}(b_{i,j},b_{\text{gt},j})\geq\tau,\,l_{i,j}=1)$ (14)
Here, $\mathcal{L}_{\text{det}}$ evaluates predictions based on spatial alignment and semantic relevance. For each candidate box $b_{i,j}$, the loss ensures geometric consistency by verifying if its IoU with the ground truth $b_{\text{gt},j}$ exceeds a threshold $\tau$. Simultaneously, the binary indicator $l_{i,j}$ confirms its validity as a positive sample. Among all valid candidates, the box with the highest IoU is selected, and its associated confidence score $Conf_{i^{*}}(p)$ is used as the detection loss. The loss focuses the adversarial signal on the most spatially aligned and semantically relevant prediction, thereby enhancing the attack effectiveness while maintaining gradient stability. Since the IoU distribution is relatively smooth and reflects spatial alignment, this approach results in a stable gradient flow during optimization. In this study, both 2D detection loss $\mathcal{L}_{\text{det}}^{\text{patch}}$ and 3D detection loss $\mathcal{L}_{\text{det}}^{\text{real}}$ are calculated using this method.
@@ 3.5.3 Patch Optimize
During adversarial sample training, the detection model is treated solely as a fixed closed-box for loss calculation and gradient propagation during patch optimization. Inspired by stochastic depth [37, 27], we employ the shakedrop regularization mechanism that randomly adjusts both the forward-output and backward-gradient flows in the residual block [45]. Specifically, in the forward pass, we combine the identity mapping with the stacked layers’ output to generate new feature representations. For an input feature $x_{\text{in}}$ and the stacked layers’ output $H(x_{\text{in}})$, the fusion rule is formulated as:
@@
where $\gamma_{1}$ is a Bernoulli random variable with $P(\gamma_{1}=1)=p_{s}$, and $\omega$ is sampled from a continuous uniform distribution $\omega\sim U(1-k,\,1+k)$. Here, $k$ is a predefined constant.
@@
where $f_{h}:[-1,1]^{d}\to\mathbb{R}^{C}$ is a pre-trained human detection model mapping the adversarial image $x^{\text{adv}}$ to class probabilities. The function $g(\cdot)$ renders the adversarial patch onto the 3D human model, ensuring its adaptation to the surfaces geometry.
The total adversarial loss for minimization is
$\mathcal{L}(f(x^{\text{adv}}(p)),y)=\mathcal{L}_{\text{total}}=\mu_{1}\cdot\mathcal{L}_{\text{det}}^{\text{patch}}+\mu_{2}\cdot\mathcal{L}_{\text{det}}^{\text{real}}+\mu_{3}\cdot\mathcal{L}_{\text{tv}}$ (17)
where the coefficients $\mu_{1}$, $\mu_{2}$, and $\mu_{3}$ are loss weights that balance the importance of 2D attacking effectiveness, 3D attacking effectiveness, and visual naturalness of the patch. Their values are set following the weighting strategy used in [27], and we maintain fixed ratios across experiments. The weighted loss formulation allows us to flexibly emphasize different optimization goals depending on specific threat models or deployment scenarios.
@@ 4.1.1 Dataset
In the Patch Adversarial Module (2D adversarial), we employ the INRIA Person dataset [46] to train and test our adversarial patch, which comprises 462 training images and 100 test images.
In the Realistic Adversarial Module (3D adversarial), we randomly selected 562 traffic scene images from the nuScenes dataset [21], captured from six camera perspectives (Front, Front Right, Front Left, Back Right, Back Left, and Back), to serve as backgrounds for 3D rendering in evaluating our adversarial patches. Among these, 450 images were captured under adequate lighting and 112 under poor lighting. The daytime and nighttime images were randomly divided into a training set of 462 images and a test set of 100 images.
@@ 4.1.2 Victim Detector
We perform adversarial training and attacks on current mainstream object detectors, including single-stage, two-stage, and transformer-based architectures (shown in Tab. 1). The models are pre-trained on MS COCO dataset [47]. In particular, we conduct closed-box attack tests on SOTA detectors (YOLO-v8, v11, v12) in the physical world for the first time. Floating Point Operations (FLOPs) quantify the computational workload required by a model during inference and serve as an indicator of the model’s complexity. In general, models with higher complexity tend to exhibit improved robustness and generalization capabilities.
@@
YOLO-v2 Single-stage 5.58 2017 [22]
YOLO-v3 193.89 2018 [23]
YOLO-v5 7.70 2020 [24]
YOLO-v8
Single-stage
8.7 2023 [42]
YOLO-v11 6.50 2024 [43]
YOLO-v12 6.50 2025 [44]
Faster-RCNN Two-stage 180.00 2016 [26]
D-DETR Transformer-based 78.00 2021 [25]
4.1.3 Evaluation Metric
In the evaluation framework, we define an attack as successful if the Intersection over Union (IoU) between the detection box and the ground-truth box is less than 0.5, or if the detection confidence score is below $0.5$. We use multiple quantitative metrics to comprehensively evaluate the effectiveness of algorithm-generated adversarial patches, as follows:
  • 1.
    Attacking Success Rate (ASR): The percentage of adversarial samples that cause the detector to fail, including false detection and missed detection. It is defined as:
    $\mathrm{ASR}=\frac{N_{\text{fail}}}{N_{\text{total}}}\times 100\%$ (18)
    where $N_{\text{fail}}$ is the number of images where the attack leads to a missed detection, and $N_{\text{total}}$ is the total number of adversarial samples.
  • 2.
    Precision: The ratio of correctly predicted positive detections to all predicted positives. Lower precision indicates that the detector is more effectively misled by the adversarial patch. It is computed as:
    $\mathrm{Precision}=\frac{TP}{TP+FP}$ (19)
    where $TP$ denotes true positives, which refer to correct detections that match ground-truth objects. $FP$ refers to false positives, meaning incorrect detections without corresponding ground-truth matches.
  • 3.
    Recall: The ratio of correctly detected ground-truth targets to the total number of ground-truth targets. Lower recall signifies that more objects are missed due to the attack. It is computed as:
    $\mathrm{Recall}=\frac{TP}{TP+FN}$ (20)
    where $FN$ indicates false negatives, which are ground-truth objects that are missed by the detector.
  • 4.
    F1-Score: The harmonic mean of precision and recall, providing a balanced measure of detection quality. It is calculated as:
    $\mathrm{F1}=2\cdot\frac{\mathrm{Precision}\cdot\mathrm{Recall}}{\mathrm{Precision}+\mathrm{Recall}}$ (21)
  • 5.
    Average Confidence (AC): The average confidence score of the detector’s predictions. A lower AC suggests that the detector is less certain about its outputs under attack. It is defined as:
    $\mathrm{AC}=\frac{1}{N}\sum_{i=1}^{N}s_{i}$ (22)
    where $s_{i}$ is the confidence score of the $i$-th detection, and $N$ is the number of valid predictions.
4.1.4 Experimental Details
To improve reproducibility and readability, we summarize the experimental setup and training hyperparameters in Tab. 2. During the digital train and evaluation, the distance from 3D human model to camera is limited in $1$ to $4$ meters. All rendered human models are at random angles and positions, and follow the principle of time-space mapping in Sec. 4.
Table 2: Experimental environment and parameters settings.
Setting Value Setting Value
GPU NVIDIA RTX 4080 (16GB) Image Resolution $416\times 416$
Patch Size (Init) $300\times 300$ Training Epochs 800
Batch Size 8 Optimizer Adam
Learning Rate 0.01 Patch Initialization Random Uniform
Loss Weights $\mu_{1}=1.0$, $\mu_{2}=1.0$, $\mu_{3}=2.5$ Random Seed 42
Camera Distance 1 to 4 meters Viewpoint Sampling Random angle and position
Non-Rigid Params $\sigma=0.8$, $\gamma=0.01$, $\rho=0.2$ Rendering Principle Time-space mapping (Sec. 4)
4.1.5 Baseline
We compared our method AdvReal with both classic and state-of-the-art (SOTA) adversarial methods. The classic methods include AdvPatch [1], AdvTshirt [14], and NatPatch [16], while the SOTA methods comprise AdvTexture [12], AdvCaT [19], and T-SEA [27]. In addition, we assess White, Gray, and Random Noise patches as control groups. For fairness, all baseline methods are evaluated using adversarial patches trained on YOLOv2 or Faster R-CNN, which are the original glass-box detectors used in their respective papers. Few of baseline methods support adversarial patch generation for more recent or structurally different detectors such as YOLOv5 or D-DETR. In contrast, our method AdvReal supports adversarial training on multiple detectors (e.g., YOLOv2, YOLOv3, YOLOv5, Faster R-CNN, D-DETR), enabling broader cross-model evaluation and generalization analysis. For fairness, the areas of all adversarial patches on both rendered and real garments are kept consistent across digital and physical experiments. With the exception of AdvCat, which is entirely camouflaged, the remaining patches and their corresponding garments are shown in Fig. 5.
4.2 Digital Adversarial Attack Experiment
4.2.1 Comparisons to SOTA Methods
@@
Method
ASR $\uparrow$
(Glass-box)
ASR $\uparrow$ (Closed-box)
ASR $\uparrow$
(Closed-box Avg)
@@
In the glass-box comparison, influenced by realistic digital world samples, the ASR of previous adversarial patches dropped considerably compared to the results reported in their original papers, revealing limited generalization under realistic deployment conditions. Moreover, the ASR of the natural-looking methods—NatPatch ($6.06\%$) and AdvCaT ($0.00\%$)—is nearly the three non-adversarial textures. In particular, the $0\%$ ASR observed for AdvCaT across multiple detectors highlights the vulnerability of camouflage-based attacks under real-world constraints. This method relies on high-frequency texture signals that are effective under controlled, close-range, high-resolution settings, but degrade significantly in realistic conditions involving longer viewing distances, varying angles, and lower resolution. Additionally, the natural appearance of the camouflage may inadvertently align with semantic cues of the “person” class, potentially reinforcing the detector’s confidence rather than suppressing it. Overall, we attribute the performance drop to training approaches that do not adequately account for real world factors such as body geometry, self-occlusion, and dynamic illumination.
Under the closed-box comparison, AdvReal achieves consistently higher attack success rates than other adversarial methods, including state-of-the-art approaches such as AdvTexture and T-SEA. Notably, when evaluated against the advanced detector YOLOv12, AdvReal achieves an ASR of $70.13\%$, dramatically outperforming T-SEA ($21.65\%$).
It is worth noting that the observed gap in adversarial performance across detectors stems from the architectural discrepancy between the black-box evaluation models and the glass-box detector used for adversarial training. Specifically, as shown in Tab.2, all adversarial patches are generated through adversarial training on YOLOv2. Since YOLOv3 shares a similar one-stage architecture with YOLOv2, the patches transfer effectively, yielding a high ASR of 77.92%. Compared to YOLO series detectors, the two-stage detectors and the transformer-based detectors adopt fundamentally different detection pipelines [53]. For example, Faster R-CNN relies on two-stage region proposal networks, while D-DETR uses attention-based token representations. The result shows that, adversarial patches against with specific architecture have significantly lower attack success rates on heterogeneous detectors [54].
To further validate AdvReal’s adversarial performance, we evaluated the ASR of various adversarial patches under different IoU thresholds. In this experiment, both YOLOv2 and Faster R-CNN are used as glass-box detectors, meaning that the adversarial patches are generated and tested under the same detector architecture to measure their optimal attack capacity. As shown in Tab. 4, AdvReal achieves higher ASRs than other attack methods when using both YOLOv2 and Faster R-CNN as detectors. Under the YOLOv2 detector, AdvReal attains an ASR of $45\%$ at an IoU threshold of 0.10.1—surpassing T-SEA’s $29.00\%$ ASR at an IoU threshold of $0.5$. With the Faster R-CNN detector, AdvReal’s ASR approaches $100\%$ at an IoU threshold of $0.5$, significantly exceeding T-SEA’s $67.10\%$. The results indicate that AdvReal can achieve nearly complete suppression of detection outputs when evaluated in a white-box setting, regardless of the detector type.
Table 4: ASRs (%) under different IoU thresholds, using YOLOv2 and Faster R-CNN as glass-box detectors and victim detectors
Method YOLOv2 (Glass-box) $\uparrow$ Faster-RCNN (Glass-box) $\uparrow$